how in the world do I find three points for the equation 3y-6=0 or 2x+5=0. I already know how to do y= 3x-2 or y=-2x+3 because the y is on the side of the problem I can do anything as long as the y is on the side the 3y-6=0 the y is not next to the equal how in the world do I solve this.
Answer by ankor@dixie-net.com(22740)

how in the world do I find three points for the equation 3y-6=0 or 2x+5=0. I already know how to do y= 3x-2 or y=-2x+3 because the y is on the side of the problem I can do anything as long as the y is on the side the 3y-6=0 the y is not next to the equal how in the world do I solve this.
"
3y - 6 = 0
Add 6 to both sides
3y = +6
Divide both side by 3
y = 2; this is plotted as a horizontal line going thru y=2, no specific value of x
:
2x + 5 = 0
subtract 5 from both sides
2x = -5
divide both sides by 2
x = -2.5
:
Plot this as a vertical line going thru x=-2.5, no specific value for y